OpenStudy (anonymous):

Identify the maximum value of the function y = -6x^2-12x-1

OpenStudy (phi):

your function is a parabola with a \( \cap \) shape. the max value is at its vertex. Can you find its vertex?

OpenStudy (igreen):

@phi is Correct. You have to find the vertex.

OpenStudy (anonymous):

max value of the function will be 5 Reason: find the 1st derivative and equate it to zero and put that value in the expression you will get the max value of the function

OpenStudy (igreen):

First start off with \(\dfrac{-b}{2a}\)

OpenStudy (igreen):

This expression will give us the x-intercept of the vertex, then we can plug it back into the equation and get our y-intercept.

OpenStudy (igreen):

Our parabola is in the form of ax^2 + bx = c -6x^2-12x-1 So in this case: a = -6 b = -12 c = -1

OpenStudy (anonymous):

@iGreen best way to find the 1st derivative rather going for the intercept

OpenStudy (igreen):

Now we can plug in our terms into the expression: \(\dfrac{-b}{2a}\) \(\dfrac{12}{2(-6)}\) \(\dfrac{12}{-12}\) -1

OpenStudy (igreen):

Which gives us -1 as the x-value for the vertex. Now we can plug in x = -1 into our parabola to get the y-intercept. \(y = -6x^2-12x-1\) \(y = -6(-1)^2-12(-1)-1\) \(y = -6(1)+12-1\) \(y = -6+12-1\) \(y = 6 - 1\) \(y = 5\) So our vertex is (-1, 5).

OpenStudy (igreen):

@LilRob

OpenStudy (anonymous):

@iGreen got it thanks man

OpenStudy (igreen):

No problem.